OpenStudy (anonymous):

Explain why \((–4x)^0 = 1\), but \(–4x^0 = –4\)

OpenStudy (turingtest):

I like messing with peoples posts I must admit :)

OpenStudy (anonymous):

anything to the 0th power is 1.

OpenStudy (cwrw238):

all numbers to the power 0 = 1

hero (hero):

Because \(-4x^0 = -4(x)^0\)

hero (hero):

The zero power is only being applied to the x

OpenStudy (cwrw238):

second one is -4 times x^0 = -4 * 1 = -4

OpenStudy (amistre64):

spose we define an exponent function notated as \(exp_b(x)\) \[exp_0(-4x)=0\] \[-4exp_0(x)=-4\]

OpenStudy (amistre64):

lol, first one would be the 1 :)

OpenStudy (turingtest):

\[(-4x)^0=(-4)^0(x)^0=1\neq-4x^0 = -4(x)^0=4(1)=4\]they are different because of the rule\[(xy)^a=x^ay^a\]in the first case there is distribution of the exponent, in the second there is not

OpenStudy (cwrw238):

(-4x)^0 - the whole term in the brackets is taken to power 0 so its = 1

OpenStudy (turingtest):

(caveat: \(x\neq0\) )

hero (hero):

@TuringTest you made a mistake

OpenStudy (turingtest):

both explanations are good however you prefer to see it

OpenStudy (cwrw238):

right - i was just wondering about x = 0 0^0 is meaningless right?

OpenStudy (turingtest):

did I? let me try to find it

hero (hero):

It might be too obvious for you to find.

OpenStudy (turingtest):

0^0 is usually regarded as an indeterminate form

OpenStudy (turingtest):

dropped the minus :/

hero (hero):

Happens to me a lot

OpenStudy (turingtest):

I meant: \[(-4x)^0=(-4)^0(x)^0=1\neq-4x^0 = -4(x)^0=-4(1)=-4\]they are different because of the rule\[(xy)^a=x^ay^a\]in the first case there is distribution of the exponent, in the second there is not better ? lol yeah it happens to us all I think, I still hate arithmetic I always say